OpenStudy (anonymous):

please help which one is true A) 1- (sin x/csc x+cot x) b) 1- (cos x /sec x +tan x)

OpenStudy (inkyvoyd):

remember than csc x=1/sin x and sec x=1/cos x; cot x=cos x/sin x; tan x=sin x/cos x

OpenStudy (inkyvoyd):

try rewriting the secants, cosecants, tangents, and cotangents as sines and cosines

OpenStudy (inkyvoyd):

Also, you've given two expressions - not exactly sure what you mean by "true" seeing as neither is an identity.

OpenStudy (debbieg):

Yeah, was just going to ask: what is there to be "true"? There is no = sign anywhere?? lol

OpenStudy (anonymous):

A) cos x B) sin x sorry

OpenStudy (inkyvoyd):

you're saying that a equals cos x or b equals sin x...?

OpenStudy (anonymous):

the question is A) cos x = 1-(sin x /csc x+cot x )\b) sin x = 1- (cos /sec x + tan x)

OpenStudy (anonymous):

is both true ?

OpenStudy (inkyvoyd):

You'll want to rewrite everything as sines and cosines.

OpenStudy (debbieg):

Is one of them definitely true? Can they both be false? @inkyvoyd were you able to show one is true? I'm not sure either one of them is true......

OpenStudy (debbieg):

If you try both for \(x=\dfrac{\pi}{4}\), neither works out. @Lindie are you sure that you have given the problem exactly as you have it?

OpenStudy (anonymous):

is it is a muli quiz

OpenStudy (debbieg):

Are these the two equations? \(\Large \cos x = 1-(\dfrac{\sin x}{\csc x}+\cot x )\) \(\Large \sin x = 1-(\dfrac{\cos x}{\sec x}+\tan x )\)

OpenStudy (debbieg):

Well, if it's a quiz, I highly doubt that you are supposed to get help answering the question.

OpenStudy (anonymous):

no it is sin x

OpenStudy (inkyvoyd):

In any case, if the two equations that @DebbieG wrote are correct, I don't see any consistency in either of them.

OpenStudy (inkyvoyd):

If it's 1-sin x/(csc x+cot x)=cos x, rewrite everything as sines and cosines and simplify the equation as you would any algebraic one.

OpenStudy (debbieg):

Yes, that makes a huuuuge difference. If you don't want to use the equation editor (which is fine), it's really really important that you carefully and EXACTLY state the COMPLETE problem. Just use ( ) carefully and it is much easier to understand. You said: cos x = 1-(sin x /csc x+cot x ) What you meant was: cos x = 1-sin x /(csc x+cot x ) ...and that one little misplaced parenthesis makes all the difference in the world. :) Like @inkyvoyd said, put everything in terms of sines and cosines, and you'll also need to use the fundamental ID along the way, as well as some factoring skills.